the cost of electricity from Electron is a flat rate of $500 plus $20/month. the cost of electricity from Vector is a flat rate of $400 plus $25/month at what point do the two companies have an equal cost


1
Expert's answer
2021-10-10T18:23:19-0400

500+20x=400+25x500+20x=400+25x

5x=1005x=100

x=1005=20x=\frac{100}{5}=20

The two companies have an equal cost at 20 months.